Pure acid is to be added to a 10% acid solution to obtain 90L of 60% solution. What amounts of each should be used?
How many liters of 100% pure acid should be used to make the solution?
How many liters of the 10% solution should be used in the mixture?

Answer by stanbon(75887)

Pure acid is to be added to a 10% acid solution to obtain 90L of 60% solution. What amounts of each should be used?
How many liters of 100% pure acid should be used to make the solution?
Equation:
1.00x + 0.10(90-x) = 0.6*90
---
100x + 10*90 - 10x = 60*90
---
90x = 50*90
x = 50 L (amt. of pure acid needed)
=============================================

How many liters of the 10% solution should be used in the mixture?
90-x = 90-50 = 40L (amt of 10% solution needed.
=================================================
Cheers,
Stan H.
============